OpenStudy (anonymous):

Rationalize the numerator:

OpenStudy (anonymous):

\[(\sqrt{x} - \sqrt{x+h}) / h \]

OpenStudy (freckles):

multiply top and bottom by conjugate of top

OpenStudy (anonymous):

I understand how to do it. You get -1 for the top, but not sure of the bottom.

OpenStudy (freckles):

\[\frac{\sqrt{x}-\sqrt{x+h}}{h} \cdot \frac{\sqrt{x}+\sqrt{x+h}}{\sqrt{x}+\sqrt{x+h}}\]

OpenStudy (freckles):

you get -1 on top before or after canceling common factors after the multiplication part

OpenStudy (freckles):

and that was a question

OpenStudy (anonymous):

I know the answer for the bottom is \[\sqrt{x} + \sqrt{x+h}\] but how do you get that with the multiplied h?

OpenStudy (freckles):

\[\frac{(x)-(x+h)}{h(\sqrt{x}+\sqrt{x+h})}\] you do understand we have -h/h=-1 right?

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

so what is the question exactly

OpenStudy (anonymous):

How does it work out for the bottom?

OpenStudy (freckles):

after using -h/h=-1 you are left with sqrt(x)+sqrt(x+h) there is nothing else to do unless you have a limit question here

OpenStudy (freckles):

sqrt(x)+sqrt(x+h) on the bottom*

OpenStudy (freckles):

\[\frac{(x)-(x+h)}{h(\sqrt{x}+\sqrt{x+h})}=\frac{-h}{h} \frac{1}{\sqrt{x}+\sqrt{x+h}}\]